How do you solve this equations: For Y

1) x-6= -1/3y
2) x-7/2=0
3) 4x-3/8y=1
4) 4x/3=3y/4+1
5) 3x/4+y/2=6

Answer by checkley77(12844)

1) x-6= -1/3y
-Y/3=X-6
-Y=3X-18
Y=-3X+18 ans.
2) x-7/2=0
NO Y TO SOLVE FOR.
3) 4x-3/8y=1
3Y/8=1-4X
3Y=8-32X
Y=(8-32X)/3 ans.
4) 4x/3=3y/4+1
3Y/4-4X/3=-1
(9Y-16X)/12=-1
9Y-16X=-12
9Y=16X-12
Y=16X/9-12/9
Y=12X/9-4/3
5) 3x/4+y/2=6
(6X+4Y)/8=6
6X+4Y=48
4Y=48-6X
Y=48/4-6Y/4
Y=12-3Y/2